OpenStudy (anonymous):

Solve the right triangle where the opposite end is a=2 and the adjacent end is b=8

OpenStudy (anonymous):

2^2 + 8^2 = c^2 4 + 64 = c^2 c= sqrt(68)

OpenStudy (anonymous):

thanks